The table shows information about a profit maximising firm.


Output 17,000 units

Price per unit $1.75

Fixed costs $ 10,000

Variable costs per unit $1.70

 

Explain whether the firm should continue production.


1
Expert's answer
2021-12-07T18:52:34-0500

Solution:

A firm should continue producing until the cost of producing the output equals the revenues obtainable from selling the output.

TC = FC + VC = 10,000 + (17,000 "\times" 1.70) = 10,000 + 28,900 = 38,900

TR = 1.75 "\times" 17,000 = 29,750

TR < TC


The firm should not continue production since total revenue is way less than its total cost, thus making huge losses.